How do I solve the system of equations by graphing y=-x+21; y=-2x+30 and find the ordered pair for this problem?
Answer by mananth(16946)

y=-x+21............1
y=-2x+30 ...........2
equation1
when x= 0 y= 21 ( 0 , 21 )
when y = 0 x= 21 ( 21 , 0 )
Equation 2
when x= 0 y= 30 ( 0 , 30 )
when y = 0 x= 15 ( 15 , 0 )

solution (9,12)